# CLIP Won’t Learn Object-Attribute Binding from Natural Data and Here is Why

### v1

@@ 5.3 Performance in the Ideal Setups for each property
Best image-incl-prob 69.85 91.76 59.76 95.67 66.31 88.30
Best caption-incl-prob 69.85 91.76 59.76 95.67 66.31 88.30
Best saliency 69.85 91.76 59.76 95.67 66.31 88.30
Best num-attrs 94.70 91.27 61.99 92.47 66.40 90.12
@@ 5.4 Recognition Performance in-distribution
(a) Multi-object image probability
(b) Multi-object caption probability
(c) Attributes per object in caption
@@
Figure 10: Impact of the data properties on the attribute recognition for in-distribution combinations. On average, it remains stable throughout, although the levels are different for the different attributes. Notably object-recognition drops for high number of attributes in the caption, likely because the model can rely more on color to distinguish samples in the batch as it becomes more ever-present.
@@ 5.5 Recognition Performance out-of-distribution
(a) Multi-object image probability
(b) Multi-object caption probability
(c) Attributes per object in caption
@@
Figure 11: Impact of the data properties on the attribute recognition for out-of-distribution combinations. It follows similar trends as in-distribution (denoted by the dotted average). However, notably the object-recognition is generally lower.

### v2

@@ 5.3 Performance in the Ideal Setups for each property
Here we look at the performance when we take the base setup of our experiments in Section˜4 and set each property separately to their ideal values. Because our base setup was already at the optimal values for three of the four properties (except Attributes-per-object-in-caption), we end up with identical and almost ideal results for most of the properties.
@@
Best Two-object-in-image-probability 69.85 91.76 59.76 95.67 66.31 88.30
Best Two-object-in-caption-probability 69.85 91.76 59.76 95.67 66.31 88.30
Best Saliency bias 69.85 91.76 59.76 95.67 66.31 88.30
Best Attributes-per-object-in-caption 94.70 91.27 61.99 92.47 66.40 90.12
@@
5.4 Effect of Individual Realistic Factors in an Ideal Setting
Here, we evaluate using the ideal setup but replacing each of the single factors with the realistic values as determined in Section˜5.2 i.e. Two-object-in-caption-probability =0.6, Two-object-in-image-probability =0.95, Saliency bias =0.9, and Attributes-per-object-in-caption with $\mathbb{E}[{n_{a}}]=0.57$.
Table 6: Evaluation using ideal setup but replacing a single factor with a realistic value. Notably, performance drops significantly when Saliency Bias is set to a realistic level, even though all other factors remain ideal.
Setup Color Scaling Fracture Rotation Swelling Thickness
Full Ideal 94.66 91.28 61.93 92.44 66.43 90.02
Ideal except Saliency bias 55.72 57.49 51.88 61.92 59.49 57.84
Ideal except Attributes-per-object-in-caption 51.39 60.28 52.46 53.69 57.61 55.64
Ideal except Two-object-in-image-probability 83.51 93.02 60.89 93.96 66.62 91.87
Ideal except Two-object-in-caption-probability 95.89 94.59 61.19 93.79 66.74 91.05
Full Realistic 50.47 54.47 52.34 50.86 53.26 50.05
@@ 5.5 Recognition Performance in-distribution
(a) Two-object-in-image-probability
(b) Two-object-in-caption-probability
(c) Attributes-per-object-in-caption
@@
Figure 10: Impact of the data properties on the attribute recognition for in-distribution combinations. On average, it remains stable throughout, although the levels are different for the different attributes. Notably object-recognition drops for high number of attributes in the caption, likely because the model can rely more on color to distinguish samples in the batch as its present with a higher probability.
@@ 5.6 Recognition Performance out-of-distribution
(a) Two-object-in-image-probability
(b) Two-object-in-caption-probability
(c) Attributes-per-object-in-caption
@@
Figure 11: Impact of the data properties on the attribute recognition for out-of-distribution combinations. It follows similar trends as in-distribution (denoted by the dotted average). However, the object-recognition is generally lower.
5.7 Influence of Data Properties on Object-Attribute Binding (No Recognition Filtering)
In Section˜4.2 we looked at the impact of the data properties under consideration on the object-attribute binding. For the results there, as detailed in Section˜3.2, we filter out samples where the attribute is not recognized for both the objects. This filtering makes sense as it factors out binding failures caused simply by failed recognitions. However, it introduces a slight variation in the test set between runs, since each evaluation uses a slightly different subset of instances. To ensure that this filtering does not significantly affect the overall trend, we also present results without recognition-based filtering. As seen in Fig.˜12 and Fig.˜13, the trend of how each of the properties influences the binding stays the same.
(a) Two-object-in-image-probability
(b) Two-object-in-caption-probability
Figure 12: Influence of (a) Two-object-in-image-probability and (b) Two-object-in-caption-probability on binding accuracy on MADMAN, evaluated without recognition filtering. The trend here is the same i.e. having more images and captions containing multiple objects improves binding.
(a) Attributes-per-object-in-caption
(b) Saliency bias
Figure 13: Influence of (a) Attributes-per-object-in-caption and (b) Saliency bias on binding accuracy on MADMAN evaluated without recognition filtering. It follows the same trend as in Fig.˜4 i.e. too few and too many Attributes-per-object-in-caption hamper the binding. And high Saliency bias is also detrimental to the binding